# GPU Cluster Scheduling for Network-Sensitive Deep Learning

### v1

@@ Abstract.
Deep learning (DL) constitutes a significant workload within public or private datacenters. Organizations commonly engage in training multiple Deep Neural Networks (DNNs) in multi-tenant GPU clusters. The primary consideration in training such models is the associated ”cost”, which correlates directly with the GPU usage duration in the cluster. Previous studies have demonstrated that communication overheads of interconnected GPU clusters could be a significant component of the overall training time and, thus, minimizing the communication overhead with intelligent scheduling of training jobs on physically close GPUs should minimize the long training times. However, state-of-the-art cluster schedulers are rather agnostic to the proximity-based job consolidation. To compound the issue, different DNN models display varied communication overheads even with similar levels of consolidation. This results in prolonged training times.
@@
In this context, we propose a novel GPU-cluster scheduler for distributed DL (DDL) workloads that enables proximity-based consolidation of GPU resources based on the DDL jobs’ sensitivities to the anticipated communication-network delays. Our scheduler consists of three major components: (i) a classical delay scheduling algorithm to facilitate job placement and consolidation; (ii) a network-sensitive job preemption strategy; and (iii) an ”auto-tuner” mechanism to optimize delay timers for effective delay scheduling. Additionally, to enable a cost-effective methodology for large-scale experiments, we develop a data-driven DDL cluster simulation platform. Employing the simulation platform we compare against several state-of-the-art alternatives on real-world workload traces to demonstrate the benefits of our design. Our scheduler can provide improvement of up to 69% in end-to-end Makespan for training all jobs compared to the prevailing consolidation-based scheduling methods, while reducing the average job completion time by up to 83% and minimizing the communication overheads by up to 98% under congested networking conditions.
copyright: none
1. INTRODUCTION
@@
Deep learning (DL) with large-scale cloud computing resources has facilitated the development of deep (very large) neural network (DNN) models, driving cutting-edge advancements in AI. These models, powering breakthroughs in domains such as autonomous vehicles (self-driving, ), multimedia processing (multimedia-DL, ) and generative AI (GAI, ), often require training hundreds of millions, if not billions or trillions, of parameters to achieve the desired accuracy. To expedite model convergence within a reasonable time-frame, parallel training paradigms based on data parallelism (pytorch_dist, ) and model parallelism (megatron, ) are employed, simultaneously utilizing multiple GPUs (or related ”neural processing” accelerators like TPUs (tpu, )) to train a single model. Furthermore, AI enterprises or groups may train hundreds or even thousands of such DNN models (philly_trace_analysis, ). In certain cases, a Machine Learning as a Service (MLaaS) (mlaas, ) provider could aggregate a substantial number of DL jobs from its diverse clientele into a single extensive GPU cluster. However, GPUs are costly, and the prolonged training times of these models lead to exorbitant training expenses.
@@
To alleviate the financial burden of DL, DNN workloads are executed on ”shared” GPU clusters, such as those offered by public cloud providers or within enterprise data centers (or private clouds). In the case of public cloud services, the cost is determined by the wallclock time spent utilizing cloud VMs, while for shared data centers, the cost is proportional to the time spent using GPU resources.
@@
Recent studies (pipedream, ; alibaba_pai, ; stash, ) have shown that communication overhead is a major contributor to prolonged training times (and increased costs) in Distributed DL (DDL). This stems from the fact that GPUs within a cluster are interconnected through various types of networks. During DNN training, collective communication algorithms such as all-reduce, scatter-gather, and others are employed to exchange tensors among GPU workers. This communication latency can vary significantly depending on the underlying network type connecting the GPUs. Generally, the more proximal (physically close) the GPUs of a job are to each other within the network, i.e., the more ”consolidated” they are, the lower the communication overhead.
@@
A DDL job can be consolidated in three typical ways: (i) placing all worker GPUs on the same physical machine; (ii) locating all worker GPUs on the same rack interconnected via Infiniband (Infiniband, ) technology; or (iii) spreading the worker GPUs across the network through inter-rack connections [^1] Hereafter referred to as the ”network.”. Each of these ”network tiers” introduces significantly different communication delays, with the least overhead on the same machine, followed by the rack, and then the network. To mitigate communication overhead, a straightforward approach involves consolidating all GPUs of a job as much as possible. However, due to multi-tenancy, when cluster managers attempt to consolidate a job on physically proximal GPUs, they often encounter delays while waiting for resources to become available, leading to increased queueing times. Furthermore, previous research (tiresias, ; stash, ) has demonstrated that DNN models exhibit varying degrees of ”network sensitivity,” implying that each DNN model would experience a different amount of increase in communication overhead as network conditions deteriorate. Thus, not all models would equally benefit from communication overhead reduction through consolidation but they would all incur the same queueing delay. Hence, it is crucial to consolidate DNN jobs based on their ”network sensitivities” i.e. relative communication overheads, in order to minimize queueing delays due to consolidation.
@@
Prior cluster scheduler Tiresias (tiresias, ) implements a consolidation strategy for DDL jobs based on the model’s ”skew.” In this strategy, jobs with high skew are consolidated into as few machines as possible, while others are left unconsolidated. This approach may lack flexibility for ”moderate consolidation,” such as placing a job on the rack for a ”moderately network-tier” sensitive model. Moderate consolidation is now more viable for communication, thanks to the presence of contemporary networking hardware that offers high quality-of-service (QoS).
@@
The contemporary Infiniband based (Infiniband, ) rack switch offers remarkably high bandwidth, reaching up to 400 Gb/s per port with the utilization of GPU RDMA (gpudirect-RDMA, ), effectively reducing rack-level communication overhead. Moreover, even modern network (Ethernet) based switches provide 800 Gb/s bandwidth per port, albeit with higher latency (nvidia-spectrum, ). Yet, DDL cluster schedulers remain oblivious to such advancements and are unable to harness the consequent potential performance improvements. To address these limitations, schedulers should be equipped with the ability to dynamically adjust consolidation based on the network sensitivity of individual jobs. This adjustment should take into account the diverse performance characteristics of different network tiers, leveraging the capabilities of modern networking hardware essential for lowering communication overhead.
@@
To achieve these goals, we propose Dally, a holistic cluster scheduler for DL job placement that accounts for the underlying system network topology, as well as the network sensitivity of each job to make intelligent placement decisions. Dally accomplishes this by employing the classical delay scheduling algorithm (delay_scheduling, ) in the context of DDL workloads, coupled with a novel network-sensitive pre-emption strategy for job placement and consolidation. Specifically, for every DDL job, Dally consults its dedicated ‘delay timers’ to decide whether to accept a resource offer at the offered network-tier or wait in anticipation of an offer with better network-tier until a delay timer has elapsed. Traditionally, the delay timers are hand-tuned and remain fixed for a given system. However, in this work, we introduce a novel ‘auto-tuner’ that dynamically evolves with the system to tune these delay timers for high-efficacy delay scheduling. With the help of these innovations, Dally is able to outperform state-of-the-art schedulers in terms of makespan, job completion time (JCT) and communication overhead.
For evaluation of Dally, we have developed a simulation platform – ArtISt-sim. The use of simulations is motivated by the fact that development of novel DL cluster schedulers is currently impeded by the exorbitant costs associated with the large-scale experimental setups needed for such endeavors. A single GPU for DL costs thousands of dollars, and replicating a real-world DL data center necessitates a substantial number of these GPUs for extended periods.
@@
ArtISt-sim has been developed by extending prior work (themis-nsdi, ) and utilizing ASTRA-sim (rashidi2020astra, ), a highly accurate DDL job simulator. Our simulation platform possesses two key capabilities: (i) it can dynamically determine network slowdowns based on the specific placement of a DDL job (its assigned GPUs) within the simulated cluster, and (ii) it can simulate a datacenter equipped with modern networking hardware (e.g., NVIDIA Quantum (nvidia-quantum, ) and Spectrum switches (nvidia-spectrum, )) which is crucial for this work. This combination of features enables ArtISt-sim to provide a realistic and accurate representation of modern DDL workloads. Moreover, to the best of our knowledge modern high speed network switches such as the ones stated before are unavailable in any public cloud for use as a test-bench. Hence, ArtISt-sim can enable researches to develop novel DDL scheduling techniques with modern networking hardware.
To summarize, in this paper, our primary contributions are outlined as follows:
@@
  • We present Dally, a novel DL cluster scheduler, crafted to reduce end-to-end Makespan, average JCT and communication overhead. Dally employs delay scheduling to achieve optimal job consolidation and placement and implements a network-sensitive preemption policy, facilitating prioritized placements of jobs with heightened sensitivity to network conditions.
@@
  • Dally also incorporates an ”auto-tuner” capable of dynamically adjusting the delay timers, based on network usage, required for delay scheduling.
@@
  • Further, we develop a high-fidelity (iteration-level) DDL cluster simulator using ASTRA-sim called ArtISt-sim, which effectively emulates a DDL data center equipped with contemporary networking hardware.
@@
  • Finally, extensive evaluation of Dally using ArtISt-sim, encompassing both batched and randomly arriving workloads of six real-world DNN models, reveals significant improvements. Dally improves makespan by over 69% compared to consolidation based policy, reduces average JCT by more than 83%, and slashes average communication overhead by 98% under congested networking conditions.
2. BACKGROUND AND RELATED WORK
2.1. Delay scheduling based on data locality
@@
Delay scheduling is a traditional scheduling algorithm used in big-data clusters that defers the scheduling of a job to the future when a suitable/preferred job placement destination (hardware) is not immediately available. Instead of promptly accepting a resource offer from any worker machine, a job undergoes a ”delay time”, anticipating the availability of a worker node with enhanced data locality in the future. Data locality refers to the location of the input data for the job, which could be on the same machine, the same rack, or somewhere accessible over the network. The farther the input data is from the job, the longer it takes for the data to be read. System administrators set the delay time, often maintaining the default values established by cluster schedulers. Traditional big-data schedulers, like YARN (yarn, ) and Spark (spark, ), typically implement delay scheduling as a ”default” strategy for better data locality. Note, deep learning jobs typically assume the availability of input data on the local machine or through fast remote storage.
2.2. Simulation Platforms
@@
ASTRA-sim (Accelerator Scaling for TRAining Simulator) is a comprehensive simulation platform facilitating parameterized descriptions of DNN model, system, and network fabric for end-to-end simulation of a DDL loop. ASTRA-sim users can configure various DDL training setups, including DNN models (number and types of layers, GEneral Matrix Multiplication (GEMM) size (gemm, ), etc), DDL parallelism (data, model, hybrid, etc.), fabric design (number of links, latency/bandwidth per link, etc.), and fabric topology (2D/3D torus, switch, ring, etc). ASTRA-sim leverages network-layer simulators, e.g., Garnet (garnet, ), Analytical (analytical, ), and ns3 (rashidi2020scalable, ), for packet-accurate simulations. An improved version of ASTRA-sim (won2023astrasim20, ) with capabilities such as arbitrary model parallelism, improved hierarchical topology parameterization and enhanced memory system modeling is now available. Also, (Akella22, ) developed a simulator using ASTRA-sim and ns3 to model various congestion control schemes.
@@
$\bullet$ System file: The system file provides parameters to the network scheduler such as topology, packet scheduling policy (e.g., LIFO, FIFO), all reduce implementation, etc.
@@
We want to emphasize however that ASTRA-sim functions as a single DDL job simulator and lacks the capability to simulate a large GPU cluster running multiple DL jobs.
2.3. Prior Work on DL cluster scheduling
A plethora of DL cluster schedulers have been developed recently with varying objectives such as JCT performance, fairness, and SLO compliance. In terms of performance, approaches like (lucid-asplos23, ; ant-man, ) focus on packing multiple DNN models onto a single GPU, boosting parallelism and decreasing overall execution time. Our work, centered around network sensitivity to enhance performance, is orthogonal to such efforts, as our proposed techniques can be readily applied to GPUs simultaneously running multiple models. However, the absence of communication characterization for such ”packed models” hinders the application of our approach. Network agnostic works such as (pollux, ; optimus, ; gandiva, ) improve performance through resource reallocation, while (horus, ) is an interference-aware scheduler which is designed for a heterogeneous GPU cluster. For fairness, (themis-nsdi, ; gandiva-fair, ) have been proposed, while (chronus, ) is an SLO compliant scheduler.
3. MOTIVATION

### v2

# Dally: A network-placement sensitive cluster scheduler for deep learning

@@ Abstract
Deep learning (DL) constitutes a significant workload within public or private cloud platforms. Organizations commonly engage in training multiple Deep Neural Networks (DNNs) in multi-tenant GPU clusters. The primary consideration in training such models is the associated “cost”, which correlates directly with the GPU usage duration in the cluster. Previous studies have demonstrated that communication latency of interconnected GPU clusters could be a significant component of the overall training time and thus, minimizing the communication latency with intelligent scheduling of training jobs on physically close GPUs should significantly reduce the long training times. However, state-of-the-art cluster schedulers are rather agnostic to the “proximity-based” job consolidation. To compound the issue, different DNN models display varied communication latency even with similar levels of consolidation. This, in turn, results in prolonged training times.
@@
Motivated by the observations above, we propose a novel GPU-cluster scheduler, Dally, for Distributed DL (DDL) workloads, that enables proximity-based consolidation of GPU resources based on the DDL jobs’ sensitivities to the anticipated communication-network delays. Dally consists of three major components: (i) a mechanism based on the classical delay scheduling algorithm to facilitate job placement and consolidation; (ii) a network-sensitive job preemption strategy; and (iii) an “auto-tuner” to optimize delay timers for effective delay scheduling. Additionally, to enable a cost-effective methodology for large-scale experiments, we develop a data-driven DDL cluster simulation platform, ArtISt-sim. Employing ArtISt-sim, we compare Dally against several state-of-the-art alternatives on real-world workload traces to demonstrate the benefits of our design. Dally can provide improvement in makespan of up to 69% (68% mean) across all training jobs, compared to the prevailing consolidation-based scheduling methods. The resultant boost in system throughput enables the reduction of average job completion time by up to 36% (26% mean) and decreases average communication latency by up to 83% (66% mean), under congested network conditions, as compared to state-of-the-art.
00footnotetext: This paper has been accepted for presentation at the IEEE Big Data 2025 conference.
I INTRODUCTION
@@
Deep learning (DL) with large-scale cloud computing resources has facilitated the development of deep (very large) neural network (DNN) models, driving cutting-edge advancements in AI. These models, powering breakthroughs in domains such as autonomous vehicles [self-driving], multimedia processing [multimedia-DL], generative AI [GAI], neuromorphic computing [nebula, skipper, gesture], text classification [Fast_Sentence_Classification] etc. often require training hundreds of millions, if not billions or trillions, of parameters to achieve the desired accuracy. To expedite model convergence within a reasonable time-frame, parallel training paradigms based on data parallelism [pytorch_dist] and model parallelism [megatron] are employed, simultaneously utilizing multiple GPUs (or related “neural processing” accelerators like TPUs [tpu]) to train a single model. Furthermore, AI enterprises or groups may train hundreds or even thousands of such DNN models [philly_trace_analysis]. In certain cases, a Machine Learning as a Service (MLaaS) [mlaas] provider could aggregate a substantial number of DL jobs from its diverse clientele into a single extensive GPU cluster. Note that MLaaS could be used both for training models from scratch or fine-tuning already-trained models. However, GPUs are costly, and the prolonged training times of these models on GPUs can easily lead to exorbitant training expenses.
@@
To alleviate the financial burden of DL, DNN workloads are executed on “shared” GPU clusters, such as those offered by public cloud providers. In such cases, the cost is determined by the wall clock time spent (makespan) utilizing cloud VMs.
@@
Recent studies [pipedream, alibaba_pai, stash] have shown that communication latency is a major contributor to prolonged training times (and thereby, increased costs) in Distributed DL (DDL). This stems from the fact that GPUs within a cluster are interconnected through various types of networks. During DNN training, collective communication algorithms such as all-reduce, scatter-gather, and others are employed to exchange tensors among GPU workers. This communication latency can vary significantly depending on the underlying network type connecting the GPUs. Generally speaking, the more proximal (physically close) the GPUs of a job are to each other within the network, i.e., the more “consolidated” they are, the lower the communication latency.
In distributed deep learning (DDL) scheduling, job placement consolidation plays a pivotal role in minimizing communication latency associated with parameter synchronization. A DDL job can be (primarily) consolidated in three typical ways: (i) placing all worker GPUs on the same (physical) machine; (ii) locating all worker GPUs on the same rack interconnected via Infiniband [Infiniband] technology; or (iii) spreading the worker GPUs across the network through inter-rack connections.
Figure 1: Single iteration training time for models consolidated on the same machine, rack, and across the network. Latency increases as the GPU workers grow physically apart.
@@
Each of these “network tiers” introduces significantly different communication latency, with the least latency being on the same machine, followed by the rack, and then the network. To illustrate this, we run a single training iteration of some popular DNNs using 8 GPUs placed across the same – (i) machine, (ii) rack and (iii) across the network, and measure their running times (shown in Fig. 1). We observe that as the GPU workers become less physically proximal on account of the underlying network placement configuration, the training latency increases. Furthermore, this increase in training latency can be attributed to the higher communication overhead incurred by the respective network. Section III-B provides further details on the experimental setup. To mitigate the communication latency, a straightforward approach involves consolidating all GPUs of a job as much as possible. However, due to multi-tenancy, when cluster managers attempt to consolidate a job on physically proximal GPUs, they often encounter delays while waiting for resources to become available, leading to increased queueing times. Furthermore, previous research [tiresias, stash] has clearly demonstrated that DNN models exhibit varying degrees of “network sensitivity,” implying that each DNN model would experience a different amount of increase in communication latency as network conditions deteriorate. Thus, not all models would equally benefit from communication latency reduction through consolidation, but they would all incur the same queueing delay. Hence, it is crucial to consolidate DNN jobs based on their “network sensitivities” i.e., relative communication latency, in order to minimize queueing delays arising from consolidation.
@@
A prior cluster scheduler, Tiresias [tiresias], implements a consolidation strategy for DDL jobs based on the “skew” of the model, which is defined as the ratio of the largest tensor size to the overall model size. In this strategy, jobs with high skew are consolidated into as few machines as possible, while others are left unconsolidated. Another prior approach [sc17-topoaware] either “packs” or “spreads” training jobs across the data-center. These approaches may lack flexibility for “moderate consolidation,” such as placing a job on the rack for a “moderately network-tier”–sensitive model. Moderate consolidation is now more viable for communication due to the availability of high speed contemporary networking hardware.
@@
The contemporary Infiniband-based [Infiniband] rack switch offers remarkably high bandwidth, reaching up to 400 Gb/s per port with the utilization of GPU RDMA [gpudirect-RDMA], effectively reducing the rack-level communication latency. Moreover, even modern network (Ethernet-based) switches provide 800 Gb/s bandwidth per port, albeit with higher link-latency  [nvidia-spectrum]. Yet, DDL cluster schedulers remain mostly oblivious to such advancements and are unable to harness the consequent potential performance improvements. To address these limitations, schedulers should be equipped with the ability to dynamically adjust consolidation based on the network sensitivity of individual jobs. This adjustment should take into account the diverse performance characteristics of different network tiers, leveraging the capabilities of modern networking hardware essential for lowering communication latency.
@@
To achieve these goals, we propose Dally, a holistic cluster scheduler for DL job placement that accounts for the underlying network topology, as well as the network sensitivity of each job to make optimal placement decisions. Dally accomplishes this by employing the classical “delay scheduling” algorithm [delay_scheduling] in the context of DDL workloads, coupled with a novel network-sensitive preemption strategy for job placement and consolidation. Specifically, for every DDL job, Dally consults its dedicated “delay timers” to decide whether to accept a resource offer at the offered network-tier or wait in anticipation of an offer with better network-tier until a delay timer has elapsed. Traditionally, the delay timers are hand-tuned and remain fixed for a given system. However, we introduce a novel “auto-tuner” that dynamically evolves with the system to tune these delay timers for high-efficacy delay scheduling. With the help of these techniques, Dally is able to outperform state-of-the-art schedulers in terms of makespan, job completion time (JCT), and communication latency.
@@
Additionally, for an accurate evaluation of Dally, we have developed a simulation platform – ArtISt-sim. The use of simulation is motivated by the fact that development of novel DL cluster schedulers is currently impeded by the exorbitant costs associated with the large-scale experimental setups needed. A single GPU for DL costs thousands of dollars, and replicating a real-world DL data center necessitates a substantial number of these GPUs for extended periods. ArtISt-sim has been developed by extending a prior work [themis-nsdi] and utilizing ASTRA-sim [rashidi2020astra], a highly accurate DDL job simulator. Our simulation platform possesses two key capabilities: (i) it can dynamically determine network slowdowns based on the specific placement of a DDL job (its assigned GPUs) within the simulated cluster, and (ii) it can simulate a datacenter equipped with modern networking hardware (e.g., NVIDIA Quantum [nvidia-quantum] and Spectrum switches [nvidia-spectrum]), which is crucial for the purposes of this work. This combination of features enables ArtISt-sim to provide a realistic and accurate representation of modern DDL workloads. Moreover, to the best of our knowledge, modern high speed network switches such as the ones stated before are unavailable in any public cloud for use as a test-bench. Hence, ArtISt-sim can enable researchers to develop novel DDL scheduling techniques with modern networking hardware, facilitating effective research in DDL cluster scheduling.
To the best of our knowledge, Dally is the first job scheduler, suitable for public cloud deployment, which introduces network-sensitive GPU consolidation policy for current and next-gen fast network hardware. This is achieved through a combination of a novel network-sensitive preemption policy and an adapted delay scheduler with an auto-tuning mechanism. Past DL cluster schedulers suffer from: (i) lack of awareness of various network tiers which prevents them from utilizing the medium-bandwidth connectivity options, such as GPU RDMA; (ii) rigid consolidation policy restricting job placement to either occupying the fewest possible machines or dispersing them across the network; (iii) inaccurate or no information about the job’s sensitivity to network stalls and; (iv) sub-optimal job prioritizing. This work also introduces a novel simulation platform for high-fidelity DL cluster simulations, ArtISt-sim.
To summarise, in this paper, our primary contributions can be outlined as follows:
@@
$\bullet$ We present Dally, a novel DL cluster scheduler, crafted through a co-design of networking hardware and underlying software, to reduce end-to-end makespan, average job completion time (JCT), and communication latency. Dally employs delay scheduling to achieve optimal job consolidation and placement and implements a network-sensitive preemption policy, facilitating prioritized placements of jobs with heightened sensitivity to network conditions.
@@
$\bullet$ Dally also incorporates an “auto-tuner” capable of dynamically adjusting the delay timers, based on network usage, required for delay scheduling.
@@
$\bullet$ Further, we develop a high-fidelity (iteration-level) DDL cluster simulator using ASTRA-sim called ArtISt-sim [artist-sim], which effectively simulates a DDL datacenter equipped with contemporary networking hardware.
@@
$\bullet$ Finally, extensive evaluation of Dally using ArtISt-sim, encompassing both batched and Poisson workloads of six real-world DNN models, reveals significant improvements. Specifically, Dally improves makespan by over 69% (68% mean) compared to state-of-the-art consolidation based policies, reduces average JCT by more than 36% (26% mean), and minimizes communication latency by up to 83% (66% mean), under congested networking conditions.
The rest of this paper is organized as follows. In Section II, we discuss the background pertinent to delay scheduling and our simulation. In Section III, we motivate our problem and our scheduler scheme is described in Section IV. The experiment methodology and evaluation is presented in Sections V and VI. Related work is in Section VII and finally, Section VIII summarizes our major observations and findings.
II BACKGROUND
II-A Delay Scheduling Based on Data Locality
@@
Delay scheduling is a traditional scheduling algorithm that defers the scheduling of a job to the future when a suitable/preferred job placement destination (hardware) is not immediately available. Instead of promptly accepting a resource offer from any worker machine, a job undergoes a “delay time”, anticipating the availability of a worker node with enhanced data locality in the future. Data locality refers to the location of the job’s input data with respect to its compute hardware, which could be on the same machine, the same rack, or somewhere accessible over the network. System administrators set the delay time, often maintaining the default values established by cluster schedulers. Traditional schedulers, like YARN [yarn] and Spark [spark], typically implement delay scheduling as a “default” strategy for better data locality.
II-B Simulation Platforms
@@
ASTRA-sim (Accelerator Scaling for TRAining Simulator) is a comprehensive simulation platform facilitating parameterized descriptions of DNN model, system, and network fabric for end-to-end simulation of a DDL loop. ASTRA-sim users can configure various DDL training setups, including DNN models (number and types of layers, GEneral Matrix Multiplication (GEMM) size [gemm], etc), DDL parallelism (data, model, hybrid, etc.), fabric design (number of links, latency/bandwidth per link, etc.), and fabric topology (2D/3D torus, switch, ring, etc). ASTRA-sim leverages existing network-layer simulators, e.g., Garnet [garnet], Analytical [analytical], and ns-3 [rashidi2020scalable], for packet-accurate simulations. An improved version of ASTRA-sim [won2023astrasim20] with capabilities such as arbitrary model parallelism, improved hierarchical topology parameterization and enhanced memory system modeling is now available. Also, [Akella22] developed a simulator using ASTRA-sim and ns-3 [rashidi2020scalable] to model various congestion control schemes.
@@
$\bullet$ System file: The system file provides parameters to the network scheduler such as topology, packet scheduling policy (e.g., LIFO, FIFO), all-reduce implementation, etc.
@@
We want to emphasize however that ASTRA-sim functions as a single DDL job simulator and lacks the capability to simulate a large GPU cluster running multiple DL jobs. Hence, in order to accurately replicate the behavior of a real-world data center, a multi job deep learning simulator is essential.
III MOTIVATION